You are trying to overhear a juicy conversation, but from your distance of 16.0 m , it sounds like only an average whisper of 30.0 dB . So you decide to move closer to give the conversation a sound level of 80.0 dB instead.

How close should you come?
Answer #1

intensity of sound is given by

b = 10 log (I/Io)

Io * 10^(b/10) = I

power = I A = Io 10^(b/10)

sincepower is to be constant

A2 Io 10^(b2/10) = A1 Io 10^(b1/10)

pi r2^2 Io 10^(80 /10) = pi r1^2 10^( 30/10)

r2 = 16 sqrt ( 10^3/10^8)

r2 = 0.05 m

=======

Goodluck